Question 2: (4 marks) Given set S is a subspace of P2, find a basis for S. s={pep, : [ Px)dx – 2p (2) = 0
Answer #1

3 ces S= {DE Pz: 5 pix) dx – 2 p(2)=0}. We know that any P (2) EP2 can be written as P(x)= ax2 + bx + C , a,b,c ER. p(x)=zaa: باایم put this value of a in p(x) we eget P(x) = ( 56 - 3c) x2+bx+c = 6(x2+x) + (-3x++1) Required basis for sis. { (- Şx2+